6. Using the following distribution of scores and the standard deviation formula of:
s = √∑(x-mean)^2/n-1

where: s = standard deviation
x = individual’s score
n = number of scores






fill in the missing values (worth 70 points):

score deviation (x – mean) (x – mean)2
student 1: 55 ______ ______ ______
student 2: 60 ______ ______ ______
student 3: 62 ______ ______ ______
student 4: 43 ______ ______ ______
student 5: 100 ______ ______ ______
student 6: 90 ______ ______ ______
------------------------------------------------------------------
∑x = ____
n = ____
mean = _____
standard deviation = _____

Answer by AnlytcPhil(1810)

6. Using the following distribution of scores and the
standard deviation formula of:

		s = √∑(x-mean)˛/(n-1) 

where: s = standard deviation
            x = individual’s score
            n = number of scores

fill in the missing values (worth 70 points):     
                                        _          _
             score     deviation = (x – x)    (x – x)˛
student 1:    55       __________________      _______
student 2:    60       __________________      _______
student 3:    62       __________________      _______
student 4:    43       __________________      _______
student 5:   100       __________________      _______
student 6:    90       __________________      _______
-------------------------------------------------------
 ∑x = ____

           n = ____
           _
    mean = x = _____

standard deviation = _____

------------------------------------------------------

Add the scores and get ∑x = 410

Count the scores and get n = 6
                               _
Divide 410 by 6 and get mean = x = 68.33333333
                                                         
Subtract this from each score to get the deviations
      _
  x - x

 55 - 68.33333333 = -13.33333333
 60 - 68.33333333 =  -8.33333333
 62 - 68.33333333 =  -6.33333333
 43 - 68.33333333 = -25.33333333
100 - 68.33333333 =  31.66666667
 90 - 68.33333333 =  21.66666667
                                 _
Square each of those to get (x - x)˛

(-13.33333333)˛ =  177.7777778
 (-8.33333333)˛ =   69.4444444
 (-6.33333333)˛ =   40.1111111
(-25.33333333)˛ =  641.7777778
 (31.66666667)˛ = 1002.7777778
 (21.66666667)˛ =  469.4444444
                       _
Add those to get ∑(x - x)˛ = 2401.333333

Divide by n-1, that is, by 6-1 which is 5

Divide by 5 and that's the variance, s˛ = 480.2666667

Take the square root of the variance to find

standard deviation = s = 21.91498726 

Edwin